# N. Deng & Y. Wang & W. Cao

### v1

@@ 3.2 Existence and uniqueness of invariant measure
$\sup_{t\geq 0}\mathbb{E}\left[\|\mathscr{O}_{t}\|^{p}_{\dot{H}^{\beta}}\right]+\sup_{t\geq 0}\mathbb{E}\left[\|\left(-A\right)^{\frac{1}{2}}\mathscr{O}_{t}\|^{p}_{W^{\gamma,2}}\right]<\infty.$ (38)
@@
$\mathbb{E}\left[\|\mathscr{O}_{t}-\mathscr{O}_{s}\|^{p}_{\dot{H}^{\beta}}\right]\leq C(t-s)^{\left(\frac{3}{8}-\frac{\beta}{4}-\epsilon\right)p},$ (39)
$\sup_{t\geq 0}\mathbb{E}\left[|o(t,x)-o(t,z)|^{p}\right]\leq C|x-z|^{p},\quad x,\ z\in\mathcal{O},$ (40)
@@
Proof 3.10.
By employing Parseval’s identity, the orthonormality of $\left\{\phi_{j}\right\}_{j\geq 0}$, Itô’s isometry, and (3), we obtain, for any $p\geq 1$ and $0\leq\beta<\frac{3}{2}$,
@@
$\leq$ $C\Big(\sum_{j=1}^{\infty}\lambda_{j}^{\beta}\Big(\big(1-e^{-\lambda_{j}^{2}(t-s)}\big)^{2}\Big\|\int_{0}^{s}e^{-\lambda_{j}^{2}(s-r)}\mathrm{d}\beta_{j}(r)\Big\|_{L^{2p}(\Omega;\mathbb{R})}^{2}+\Big\|\int_{s}^{t}e^{-\lambda_{j}^{2}(t-r)}\mathrm{d}\beta_{j}(r)\Big\|_{L^{2p}(\Omega;\mathbb{R})}^{2}\Big)\Big)^{\frac{1}{2}}$
@@
$\sup_{t\geq 0}\mathbb{E}\left[\|\mathscr{O}_{t}\|^{p}_{\dot{H}^{\beta}}\right]<\infty.$ (41)
@@
$\left\|\partial_{x}o(t,x)-\partial_{y}o(t,y)\right\|^{2}_{L^{p}(\Omega;\mathbb{R})}\leq$ $C\sum_{j=1}^{\infty}\int_{0}^{t}\lambda_{j}e^{-2\lambda_{j}^{2}(t-s)}|\psi_{j}(x)-\psi_{j}(y)|^{2}\mathrm{d}s\leq C\sum_{j=1}^{\infty}\lambda_{j}^{-1}|\psi_{j}(x)-\psi_{j}(y)|^{2}$ (42)
@@
$\left\|\left(\int_{\mathcal{O}}\int_{\mathcal{O}}\frac{|\partial_{x}o(t,x)-\partial_{y}o(t,y)|^{2}}{|x-y|^{1+2\gamma}}\mathrm{d}x\mathrm{d}y\right)^{\frac{1}{2}}\right\|_{L^{p}(\Omega;\mathbb{R})}\leq$ $\left(\int_{\mathcal{O}}\int_{\mathcal{O}}\frac{\left\|\partial_{x}o(t,x)-\partial_{y}o(t,y)\right\|^{2}_{L^{p}(\Omega;\mathbb{R})}}{|x-y|^{1+2\gamma}}\mathrm{d}x\mathrm{d}y\right)^{\frac{1}{2}}$ (43)
@@
Combining (41), (43), and the following identy
@@
we obtain (38). In addition, (40) follows from an argument analogous to that of (42). This completes the proof.
@@
Suppose that Assumptions 22 hold. Then for any $p\geq 1$, it holds that
$\sup_{t\in[0,\infty)}\mathbb{E}\left[\|u(t,\cdot)\|^{p}_{L^{\infty}}\right]<\infty.$ (44)
Proof 3.12.
It follows from (16), the Sobolev embedding theorem, and Lemma 3.9 that
$\sup_{t\geq 0}\mathbb{E}\left[\|\mathscr{O}_{t}\|^{q}_{L^{\infty}}\right]+\sup_{t\geq 0}\mathbb{E}\left[\|(-A)^{\frac{1}{2}}\mathscr{O}_{t}\|^{q}_{L^{4}}\right]$ (45)
@@
Combining (14), (45) with Proposition 3.2 yields
@@
$\sup_{t\geq 0}\mathbb{E}\left[\|u(t,\cdot)\|^{p}_{L^{\infty}}\right]=\sup_{t\geq 0}\mathbb{E}\left[\|U_{t}\|^{p}_{L^{\infty}}\right]\leq C\sup_{t\geq 0}\mathbb{E}\left[\|U_{t}\|^{p}_{W^{1,2}}\right]<\infty.$ (46)
The proof is completed.
By using Theorem 3.11 and Corollary 3.7, we present the existence of invariant measure for the SCHE (1) in the following theorem. We recall that $H_{\alpha}:=\left\{v\in\dot{H}:\frac{1}{\pi}\int_{\mathcal{O}}v\mathrm{d}x=\alpha\right\}$. The following analysis is valid for any fixed $\alpha$.
@@
Suppose that Assumptions 2-2 hold, then there exists an invariant measure $\pi$ for the SCHE (1) on $H_{\alpha}$.
Proof 3.14.
@@
This implies that the family $\left\{\hat{\pi}_{t}\right\}_{t\in[0,\infty)}$ is tight. Consequently, by Corollary 3.7 and the Krylov–Bogoliubov Theorem (Hong & Wang, 2019, Theorem 1.2), the mild solution (9) admits an invariant measure $\pi$ on $H_{\alpha}$.
Da Prato & Debussche (1996) has established the uniqueness of the invariant measures on $\dot{H}^{-1}_{\alpha}$. Next we aim to extend this result to $H_{\alpha}$. To this end, we first establish the following lemma to explain the relationship between the uniqueness of the invariant measure on different spaces.
@@
Suppose that Assumptions 2-2 hold, then the invariant measure $\pi$ for the SCHE (1) is unique on $H_{\alpha}$.
Proof 3.18.
According to Da Prato & Debussche (1996, Theorem 3.3), the mild solution given by (9) admits an unique invariant measure $\pi$ on $\dot{H}^{-1}_{\alpha}$. Since $H_{\alpha}$ is a complete subspace of $\dot{H}^{-1}_{\alpha}$ that is continuously embedded into $\dot{H}^{-1}_{\alpha}$ and $u(t,\cdot)\in H_{\alpha}$ for all $t\geq 0$, it follows from Lemma 3.15 and Theorem 3.13 that there exists a unique invariant measure $\pi$ for the SCHE (1) on $H_{\alpha}$.
@@
4 Fully discrete scheme and uniform strong convergence
In this section, following the methods developed in Deng& Cao (2025), we introduce a fully discrete scheme for original equation. We use a finite difference method for the spatial discretization and a tamed exponential Euler method for the temporal discretization. We establish the uniform strong convergence of our proposed scheme.
@@ 4 Fully discrete scheme and uniform strong convergence
See Sun et al. (2023, Chapter 9) for details. Considering Eq. (1) at the point $(t,x_{i})$, we approximate $\Delta$ and $\Delta^{2}$ by $\delta_{h}^{2}$ and $\delta_{h}^{4}$. $W(t,x_{i})$ is approximated by $\sum_{j=1}^{N-1}\phi_{j}(x_{i})\beta_{j}(t)$. In this way, we establish a scheme given by
$\mathrm{d}u^{N}\left(t,x_{i}\right)+\delta_{h}^{4}u^{N}\left(t,x_{i}\right)\mathrm{d}t=\delta_{h}^{2}f\left(u^{N}\left(t,x_{i}\right)\right)\mathrm{d}t+\sigma\sum_{j=1}^{N-1}\phi_{j}(x_{i})\mathrm{d}\beta_{j}(t).$ (47)
Note that (47) is defined only at the grid points. In order to construct its spatial continuous extension, it is necessary to introduce two operators. We introduce the operator $\mathcal{S}_{N}:\ H\to l_{N}^{2}$ defined by
@@
where $\left\{\phi_{j}\right\}_{j=0}^{\infty}$ is an orthonormal basis of $\dot{H}$. Then it follows from (47) that
$\mathrm{d}U_{t}^{N}+A_{N}^{2}U_{t}^{N}\mathrm{d}t=A_{N}F_{N}(U_{t}^{N})\mathrm{d}t+\sigma\sum_{j=1}^{N-1}\phi_{N,j}\mathrm{d}\beta_{j}(t),\quad t>0,\$(48a)
$U_{0}^{N}=u^{N}_{0},\$(48b)
@@
$\|\vec{u}\|_{l_{N}^{\infty}}\leq C\|\vec{u}\|_{w^{1,2}_{N}},$ (49)
$\|\vec{w}\|_{w^{1,2}_{N}}\leq C\|\vec{u}\|_{w^{1,2}_{N}}\|\vec{v}\|_{w^{1,2}_{N}},$ (50)
$\|\vec{u}\|_{l_{N}^{6}}\leq C\Big(\left\|A_{N}\vec{u}\right\|_{l_{N}^{2}}^{\frac{1}{6}}\|\vec{u}\|_{l_{N}^{2}}^{\frac{5}{6}}+\|\vec{u}\|_{l_{N}^{2}}\Big),$ (51)
$\left\|(-A_{N})^{\gamma}e^{-A_{N}^{2}t}\vec{u}\right\|_{l_{N}^{2}}\leq C\Big(t^{-\frac{\gamma}{2}}e^{-\frac{\lambda_{N,1}^{2}}{2}t}\left\|\vec{u}\right\|_{l_{N}^{2}}+g(\gamma)\big|\left\langle\vec{u},\phi_{N,0}\right\rangle_{l_{N}^{2}}\big|\Big),\ t>0,\$(52)
$\left\|(-A_{N})^{\gamma_{1}}e^{-A_{N}^{2}t}\left(I-e^{-A_{N}^{2}s}\right)\vec{u}\right\|_{l_{N}^{2}}\leq C\Big(t^{-\frac{\gamma_{1}-\gamma_{2}+2\gamma_{3}}{2}}s^{\gamma_{3}}e^{-\frac{\lambda_{N,1}^{2}}{2}t}\left\|(-A_{N})^{\gamma_{2}}\vec{u}\right\|_{l_{N}^{2}}+g(\gamma_{1})\big|\left\langle\hat{u},\phi_{N,0}\right\rangle_{l_{N}^{2}}\big|\Big),\ s,t>0,$ (53)
@@
Inequalities (49)-(51) can be found in Theorem 2.1 of Deng& Cao (2025) and the proof of inequalities (52)-(53) is similar to that of (4)-(5). Therefore, the proof of Lemma 4.1 is omitted.
@@
$U^{\tau,N}_{t}=$ $e^{-A_{N}^{2}t}u_{0}^{N}+\int_{0}^{t}A_{N}e^{-A_{N}^{2}(t-\lfloor s\rfloor)}\tilde{F}_{N}(U^{\tau,N}_{\lfloor s\rfloor})\mathrm{d}s+\sigma\sum_{j=1}^{N-1}\int_{0}^{t}e^{-A_{N}^{2}(t-\lfloor s\rfloor)}\phi_{N,j}\mathrm{d}\beta_{j}(s),$ (54)
@@
We are aware that some recent studies (e.g., Deng& Cao (2025), Qi & Wang (2020), Qi et al. (2024)) have proved strong convergence of fully discrete schemes for the SCHE on finite-time intervals. In our work, we aim to extend these convergence results from finite intervals to the infinite-time case. Inspired by Wang & Cao (2024), we enhance the tamed term in our numerical scheme by replacing $1/\big(1+\tau\|F_{N}(\cdot)\|_{l^{2}_{N}}\big)$ with $1/\big(1+\tau\|\cdot\|_{w^{1,2}_{N}}^{12}\big)$. The design of novel tamed strategies has attracted increasing attention, as exemplified by recent works such as Jiang & Wang (2025) and Liu & Shen (2025). In addition, the fully discrete scheme (54) can be rewritten as
$U^{\tau,N}_{t_{m+1}}=$ $e^{-A_{N}^{2}\tau}U^{\tau,N}_{t_{m}}+\tau A_{N}e^{-A_{N}^{2}\tau}\tilde{F}_{N}(U^{\tau,N}_{t_{m}})+\sigma e^{-A_{N}^{2}\tau}\Delta\beta_{m},\quad m\in\mathbb{N},$ (55)
@@
In order to introduce the spatially continuous form of (54), we define
@@
$u^{\tau,N}(t,x)=P_{N}U^{\tau,N}_{t}(x)=$ $\int_{\mathcal{O}}G_{t}^{N}(x,y)u_{0}\left(\kappa_{N}(y)\right)\mathrm{d}y+\int_{0}^{t}\int_{\mathcal{O}}\frac{\Delta_{N}G_{t-\lfloor s\rfloor}^{N}(x,y)f\left(u^{\tau,N}\left(\lfloor s\rfloor,\kappa_{N}(y)\right)\right)}{1+\tau\left\|U^{\tau,N}_{\lfloor s\rfloor}\right\|_{w^{1,2}_{N}}^{12}}\mathrm{d}y\mathrm{d}s$ (56)
@@
$\sup_{x\in\mathcal{O}}\int_{\mathcal{O}}\left|\Delta_{N}G_{t}^{N}(x,y)-\Delta G_{t}(x,y)\right|\mathrm{d}y\leq Ce^{-\lambda_{N,1}t}t^{-\frac{2\alpha_{1}+5}{8}}N^{-\alpha_{1}},\quad t>0.$ (57)
4.1 Uniform moment boundedness of the numerical solution
@@
$\sup_{t\geq 0}\mathbb{E}\left[\big\|(-A_{N})^{\frac{\beta}{2}}\mathscr{O}^{N}_{t}\big\|^{p}_{l^{2}_{N}}\right]+\sup_{t\geq 0}\mathbb{E}\left[\big\|\left(-A_{N}\right)^{\frac{1}{2}}\mathscr{O}^{N}_{t}\big\|^{p}_{w^{\gamma,2}_{N}}\right]<\infty,$ (58)
$\sup_{t\geq 0}\mathbb{E}\left[\big\|(-A_{N})^{\frac{\beta}{2}}\mathscr{O}^{\tau,N}_{t}\big\|^{p}_{l^{2}_{N}}\right]+\sup_{t\geq 0}\mathbb{E}\left[\big\|\left(-A_{N}\right)^{\frac{1}{2}}\mathscr{O}^{\tau,N}_{t}\big\|^{p}_{w^{\gamma,2}_{N}}\right]<\infty.$ (59)
@@
$\mathbb{E}\left[\big\|(-A_{N})^{\frac{\beta}{2}}\left(\mathscr{O}^{\tau,N}_{t}-\mathscr{O}^{\tau,N}_{s}\right)\big\|^{p}_{l^{2}_{N}}\right]+\mathbb{E}\left[\big\|(-A_{N})^{\frac{\beta}{2}}\left(\mathscr{O}^{\tau,N}_{t}-\mathscr{O}^{\tau,N}_{s}\right)\big\|^{p}_{l^{2}_{N}}\right]\leq C(t-s)^{\left(\frac{3}{8}-\frac{\beta}{4}-\epsilon\right)p},$ (60)
$\sup_{t\geq 0}\mathbb{E}\left[|o^{N}(t,x)-o^{N}(t,z)|^{p}\right]+\sup_{{t\geq 0}}\mathbb{E}\left[|o^{\tau,N}(t,x)-o^{\tau,N}(t,z)|^{p}\right]\leq C|x-z|^{p},\quad x,\ z\in\mathcal{O},$ (61)
@@
Suppose that Assumptions 2-2 hold. Then for any $p\geq 2$, it holds that
$\sup_{\tau\in(0,1),N\in\mathbb{N}}\sup_{i\geq 0}\mathbb{E}\left[\left\|u^{\tau,N}(t_{m},\cdot)\right\|_{l^{\infty}_{N}}^{p}\right]<\infty\quad\text{and}\quad\sup_{\tau\in(0,1),N\in\mathbb{N}}\sup_{i\geq 0}\mathbb{E}\left[\left\|U_{t_{i}}^{\tau,N}\right\|_{w^{1,2}_{N}}^{p}\right]<\infty.$ (62)
Proof 4.6.

### v3

@@ 3.  Invariant measure for the stochastic Cahn–Hilliard equation
$\sup_{t\geq 0}\mathbb{E}\left[\|\mathscr{O}_{t}\|^{p}_{\dot{H}^{\beta}}\right]+\sup_{t\geq 0}\mathbb{E}\left[\|\left(-A\right)^{\frac{1}{2}}\mathscr{O}_{t}\|^{p}_{W^{\gamma,2}}\right]<\infty.$ (3.30)
@@
$\mathbb{E}\left[\|\mathscr{O}_{t}-\mathscr{O}_{s}\|^{p}_{\dot{H}^{\beta}}\right]\leq C(t-s)^{\left(\frac{3}{8}-\frac{\beta}{4}-\epsilon\right)p},$ (3.31)
$\sup_{t\geq 0}\mathbb{E}\left[|o(t,x)-o(t,z)|^{p}\right]\leq C|x-z|^{p},\quad x,\ z\in\mathcal{O},$ (3.32)
@@
Proof.
By employing Parseval’s identity, the orthonormality of $\left\{\phi_{j}\right\}_{j\geq 0}$, Itô’s isometry, and (2.2), we obtain, for any $p\geq 1$ and $0\leq\beta<\frac{3}{2}$,
@@
$\leq$ $C\Big(\sum_{j=1}^{\infty}\lambda_{j}^{\beta}\Big(\big(1-e^{-\lambda_{j}^{2}(t-s)}\big)^{2}\Big\|\int_{0}^{s}e^{-\lambda_{j}^{2}(s-r)}\mathrm{d}\beta_{j}(r)\Big\|_{L^{2p}(\Omega;\mathbb{R})}^{2}$
$\qquad\qquad\qquad+\Big\|\int_{s}^{t}e^{-\lambda_{j}^{2}(t-r)}\mathrm{d}\beta_{j}(r)\Big\|_{L^{2p}(\Omega;\mathbb{R})}^{2}\Big)\Big)^{\frac{1}{2}}$
@@
$\sup_{t\geq 0}\mathbb{E}\left[\|\mathscr{O}_{t}\|^{p}_{\dot{H}^{\beta}}\right]<\infty.$ (3.33)
@@
$\left\|\partial_{x}o(t,x)-\partial_{y}o(t,y)\right\|^{2}_{L^{p}(\Omega;\mathbb{R})}\leq$ $C\sum_{j=1}^{\infty}\int_{0}^{t}\lambda_{j}e^{-2\lambda_{j}^{2}(t-s)}|\psi_{j}(x)-\psi_{j}(y)|^{2}\mathrm{d}s\leq C\sum_{j=1}^{\infty}\lambda_{j}^{-1}|\psi_{j}(x)-\psi_{j}(y)|^{2}$ (3.34)
@@
$\left\|\left(\int_{\mathcal{O}}\int_{\mathcal{O}}\frac{|\partial_{x}o(t,x)-\partial_{y}o(t,y)|^{2}}{|x-y|^{1+2\gamma}}\mathrm{d}x\mathrm{d}y\right)^{\frac{1}{2}}\right\|_{L^{p}(\Omega;\mathbb{R})}\leq$ $\left(\int_{\mathcal{O}}\int_{\mathcal{O}}\frac{\left\|\partial_{x}o(t,x)-\partial_{y}o(t,y)\right\|^{2}_{L^{p}(\Omega;\mathbb{R})}}{|x-y|^{1+2\gamma}}\mathrm{d}x\mathrm{d}y\right)^{\frac{1}{2}}$ (3.35)
@@
Combining (3.33), (3.35), and the following identy
@@
we obtain (3.30). In addition, (3.32) follows from an argument analogous to that of (3.34). This completes the proof. ∎
@@
Suppose that Assumptions 2.22.4 hold. Then for any $p\geq 1$, it holds that
$\sup_{t\in[0,\infty)}\mathbb{E}\left[\|u(t,\cdot)\|^{p}_{L^{\infty}}\right]<\infty.$ (3.36)
Proof.
It follows from (3.8), the Sobolev embedding theorem, and Lemma 3.6 that
$\sup_{t\geq 0}\mathbb{E}\left[\|\mathscr{O}_{t}\|^{q}_{L^{\infty}}\right]+\sup_{t\geq 0}\mathbb{E}\left[\|(-A)^{\frac{1}{2}}\mathscr{O}_{t}\|^{q}_{L^{4}}\right]$ (3.37)
@@
Combining (3.6), (3.37) with Proposition 3.2 yields
@@
$\sup_{t\geq 0}\mathbb{E}\left[\|u(t,\cdot)\|^{p}_{L^{\infty}}\right]=\sup_{t\geq 0}\mathbb{E}\left[\|U_{t}\|^{p}_{L^{\infty}}\right]\leq C\sup_{t\geq 0}\mathbb{E}\left[\|U_{t}\|^{p}_{W^{1,2}}\right]<\infty.$ (3.38)
The proof is completed. ∎
By using Theorem 3.1 and Corollary 3.5, we present the existence of invariant measure for the SCHE (1.1) in the following theorem. We recall that $H_{\alpha}:=\left\{v\in\dot{H}:\frac{1}{\pi}\int_{\mathcal{O}}v\mathrm{d}x=\alpha\right\}$. The following analysis is valid for any fixed $\alpha$.
@@
Suppose that Assumptions 2.2-2.4 hold, then there exists an invariant measure $\pi$ for the SCHE (1.1) on $H_{\alpha}$.
Proof.
@@
This implies that the family $\left\{\hat{\pi}_{t}\right\}_{t\in[0,\infty)}$ is tight. Consequently, by Corollary 3.5 and the Krylov–Bogoliubov Theorem [24, Theorem 1.2], the mild solution (3.1) admits an invariant measure $\pi$ on $H_{\alpha}$. ∎
[15] has established the uniqueness of the invariant measures on $\dot{H}^{-1}_{\alpha}$. Next we aim to extend this result to $H_{\alpha}$. To this end, we first establish the following lemma to explain the relationship between the uniqueness of the invariant measure on different spaces.
@@
Suppose that Assumptions 2.2-2.4 hold, then the invariant measure $\pi$ for the SCHE (1.1) is unique on $H_{\alpha}$.
Proof.
According to [15, Theorem 3.3], the mild solution given by (3.1) admits an unique invariant measure $\pi$ on $\dot{H}^{-1}_{\alpha}$. Since $H_{\alpha}$ is a complete subspace of $\dot{H}^{-1}_{\alpha}$ that is continuously embedded into $\dot{H}^{-1}_{\alpha}$ and $u(t,\cdot)\in H_{\alpha}$ for all $t\geq 0$, it follows from Lemma 3.7 and Theorem 3.2 that there exists a unique invariant measure $\pi$ for the SCHE (1.1) on $H_{\alpha}$. ∎
@@
4.  Fully discrete scheme and uniform strong convergence
In this section, following the methods developed in [19], we introduce a fully discrete scheme for original equation. We use a finite difference method for the spatial discretization and a tamed exponential Euler method for the temporal discretization. We establish the uniform strong convergence of our proposed scheme.
@@ 4.  Fully discrete scheme and uniform strong convergence
See [37, Chapter 9] for details. Considering Eq. (1.1) at the point $(t,x_{i})$, we approximate $\Delta$ and $\Delta^{2}$ by $\delta_{h}^{2}$ and $\delta_{h}^{4}$. $W(t,x_{i})$ is approximated by $\sum_{j=1}^{N-1}\phi_{j}(x_{i})\beta_{j}(t)$. In this way, we establish a scheme given by
$\mathrm{d}u^{N}\left(t,x_{i}\right)+\delta_{h}^{4}u^{N}\left(t,x_{i}\right)\mathrm{d}t=\delta_{h}^{2}f\left(u^{N}\left(t,x_{i}\right)\right)\mathrm{d}t+\sigma\sum_{j=1}^{N-1}\phi_{j}(x_{i})\mathrm{d}\beta_{j}(t).$ (4.1)
Note that (4.1) is defined only at the grid points. In order to construct its spatial continuous extension, it is necessary to introduce two operators. We introduce the operator $\mathcal{S}_{N}:\ H\to l_{N}^{2}$ defined by
@@
where $\left\{\phi_{j}\right\}_{j=0}^{\infty}$ is an orthonormal basis of $\dot{H}$. Then it follows from (4.1) that
$\mathrm{d}U_{t}^{N}+A_{N}^{2}U_{t}^{N}\mathrm{d}t=A_{N}F_{N}(U_{t}^{N})\mathrm{d}t+\sigma\sum_{j=1}^{N-1}\phi_{N,j}\mathrm{d}\beta_{j}(t),\quad t>0,\$(4.2a)
$U_{0}^{N}=u^{N}_{0},\$(4.2b)
@@
$\|\vec{u}\|_{l_{N}^{\infty}}\leq C\|\vec{u}\|_{w^{1,2}_{N}},$ (4.3)
$\|\vec{w}\|_{w^{1,2}_{N}}\leq C\|\vec{u}\|_{w^{1,2}_{N}}\|\vec{v}\|_{w^{1,2}_{N}},$ (4.4)
$\|\vec{u}\|_{l_{N}^{6}}\leq C\Big(\left\|A_{N}\vec{u}\right\|_{l_{N}^{2}}^{\frac{1}{6}}\|\vec{u}\|_{l_{N}^{2}}^{\frac{5}{6}}+\|\vec{u}\|_{l_{N}^{2}}\Big),$ (4.5)
$\left\|(-A_{N})^{\gamma}e^{-A_{N}^{2}t}\vec{u}\right\|_{l_{N}^{2}}\leq C\Big(t^{-\frac{\gamma}{2}}e^{-\frac{\lambda_{N,1}^{2}}{2}t}\left\|\vec{u}\right\|_{l_{N}^{2}}+g(\gamma)\big|\left\langle\vec{u},\phi_{N,0}\right\rangle_{l_{N}^{2}}\big|\Big),\ t>0,\$(4.6)
$\left\|(-A_{N})^{\gamma_{1}}e^{-A_{N}^{2}t}\left(I-e^{-A_{N}^{2}s}\right)\vec{u}\right\|_{l_{N}^{2}}\leq C\Big(t^{-\frac{\gamma_{1}-\gamma_{2}+2\gamma_{3}}{2}}s^{\gamma_{3}}e^{-\frac{\lambda_{N,1}^{2}}{2}t}\left\|(-A_{N})^{\gamma_{2}}\vec{u}\right\|_{l_{N}^{2}}+g(\gamma_{1})\big|\left\langle\hat{u},\phi_{N,0}\right\rangle_{l_{N}^{2}}\big|\Big),$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad s,t>0,$ (4.7)
@@
Inequalities (4.3)-(4.5) can be found in Theorem 2.1 of [19] and the proof of inequalities (4.6)-(4.7) is similar to that of (2.3)-(2.4). Therefore, the proof of Lemma 4.1 is omitted.
@@
$U^{\tau,N}_{t}=$ $e^{-A_{N}^{2}t}u_{0}^{N}+\int_{0}^{t}A_{N}e^{-A_{N}^{2}(t-\lfloor s\rfloor)}\tilde{F}_{N}(U^{\tau,N}_{\lfloor s\rfloor})\mathrm{d}s+\sigma\sum_{j=1}^{N-1}\int_{0}^{t}e^{-A_{N}^{2}(t-\lfloor s\rfloor)}\phi_{N,j}\mathrm{d}\beta_{j}(s),$ (4.8)
@@
We are aware that some recent studies (e.g., [19], [33], [32]) have proved strong convergence of fully discrete schemes for the SCHE on finite-time intervals. In our work, we aim to extend these convergence results from finite intervals to the infinite-time case. Inspired by [35], we enhance the tamed term in our numerical scheme by replacing $1/\big(1+\tau\|F_{N}(\cdot)\|_{l^{2}_{N}}\big)$ with $1/\big(1+\tau\|\cdot\|_{w^{1,2}_{N}}^{12}\big)$. The design of novel tamed strategies has attracted increasing attention, as exemplified by recent works such as [25] and [28]. In addition, the fully discrete scheme (4.8) can be rewritten as
$U^{\tau,N}_{t_{m+1}}=$ $e^{-A_{N}^{2}\tau}U^{\tau,N}_{t_{m}}+\tau A_{N}e^{-A_{N}^{2}\tau}\tilde{F}_{N}(U^{\tau,N}_{t_{m}})+\sigma e^{-A_{N}^{2}\tau}\Delta\beta_{m},\quad m\in\mathbb{N},$ (4.9)
@@
In order to introduce the spatially continuous form of (4.8), we define
@@
$u^{\tau,N}(t,x)=P_{N}U^{\tau,N}_{t}(x)=$ $\int_{\mathcal{O}}G_{t}^{N}(x,y)u_{0}\left(\kappa_{N}(y)\right)\mathrm{d}y+\int_{0}^{t}\int_{\mathcal{O}}\frac{\Delta_{N}G_{t-\lfloor s\rfloor}^{N}(x,y)f\left(u^{\tau,N}\left(\lfloor s\rfloor,\kappa_{N}(y)\right)\right)}{1+\tau\left\|U^{\tau,N}_{\lfloor s\rfloor}\right\|_{w^{1,2}_{N}}^{12}}\mathrm{d}y\mathrm{d}s$ (4.10)
@@
$\sup_{x\in\mathcal{O}}\int_{\mathcal{O}}\left|\Delta_{N}G_{t}^{N}(x,y)-\Delta G_{t}(x,y)\right|\mathrm{d}y\leq Ce^{-\lambda_{N,1}t}t^{-\frac{2\alpha_{1}+5}{8}}N^{-\alpha_{1}},\quad t>0.$ (4.11)
4.Uniform moment boundedness of the numerical solution
@@
$\sup_{t\geq 0}\mathbb{E}\left[\big\|(-A_{N})^{\frac{\beta}{2}}\mathscr{O}^{N}_{t}\big\|^{p}_{l^{2}_{N}}\right]+\sup_{t\geq 0}\mathbb{E}\left[\big\|\left(-A_{N}\right)^{\frac{1}{2}}\mathscr{O}^{N}_{t}\big\|^{p}_{w^{\gamma,2}_{N}}\right]<\infty,$ (4.12)
$\sup_{t\geq 0}\mathbb{E}\left[\big\|(-A_{N})^{\frac{\beta}{2}}\mathscr{O}^{\tau,N}_{t}\big\|^{p}_{l^{2}_{N}}\right]+\sup_{t\geq 0}\mathbb{E}\left[\big\|\left(-A_{N}\right)^{\frac{1}{2}}\mathscr{O}^{\tau,N}_{t}\big\|^{p}_{w^{\gamma,2}_{N}}\right]<\infty.$ (4.13)
@@
$\mathbb{E}\left[\big\|(-A_{N})^{\frac{\beta}{2}}\left(\mathscr{O}^{\tau,N}_{t}-\mathscr{O}^{\tau,N}_{s}\right)\big\|^{p}_{l^{2}_{N}}\right]+\mathbb{E}\left[\big\|(-A_{N})^{\frac{\beta}{2}}\left(\mathscr{O}^{\tau,N}_{t}-\mathscr{O}^{\tau,N}_{s}\right)\big\|^{p}_{l^{2}_{N}}\right]\leq C(t-s)^{\left(\frac{3}{8}-\frac{\beta}{4}-\epsilon\right)p},$ (4.14)
$\sup_{t\geq 0}\mathbb{E}\left[|o^{N}(t,x)-o^{N}(t,z)|^{p}\right]+\sup_{{t\geq 0}}\mathbb{E}\left[|o^{\tau,N}(t,x)-o^{\tau,N}(t,z)|^{p}\right]\leq C|x-z|^{p},\quad x,\ z\in\mathcal{O},$ (4.15)
@@
Suppose that Assumptions 2.2-2.4 hold. Then for any $p\geq 2$, it holds that
$\sup_{\tau\in(0,1),N\in\mathbb{N}}\sup_{i\geq 0}\mathbb{E}\left[\left\|u^{\tau,N}(t_{m},\cdot)\right\|_{l^{\infty}_{N}}^{p}\right]<\infty\quad\text{and}\quad\sup_{\tau\in(0,1),N\in\mathbb{N}}\sup_{i\geq 0}\mathbb{E}\left[\left\|U_{t_{i}}^{\tau,N}\right\|_{w^{1,2}_{N}}^{p}\right]<\infty.$ (4.16)
Proof.